The sum of the ages of a father and her daughter is 44 years. The father's age is 4 years more than 3 times the daughter's age. Find the ages of both.
Answer by elima(1433)

The sum of the ages of a father and her daughter is 44 years. The father's age is years more than 3 times the daughter's age. Find the ages of both.
Let x be the daughter's age
The fathers age is:
4+3x
So the eqaution would be:
x+(4+3x)=44
We combine like terms:
x+3x+4=44
4x+4=44
We subtract 4 from both sides:
4x=40
We divide by 4 from both sides:
x=10 - daughters age
10+(4+3(10)=44
10+4+30=44
So if the both add to 44 and the daughters age is 10 then the fathers age must be 34.